Oxidation number of Fe in Fe0.94O is-
a.200    b.200/94     c.94/200

Grade:

1 Answers

jyoti bhatia
5 years ago
200/94 is correct option.
Usually, Fe in oxides will be in either +2 oxidation state or in +3 oxidation state. Fe0.94O is a non-stoichiometric compound. A non-stoichiometric compound is that in which the cation/anion ratio is not simple.
 
Let us suppose that the oxidation number of iron is x.

We know that the oxidation number of O is −2.

Therefore, oxidation number of iron can be calculated as:
0.94x − 2 = 0
 
x = 2/0.94 = 200/94 = 2.12
 
Therefore, the oxidation number of Fe in Fe0.94O is = 200/94 = 2.12